OpenStudy (anonymous):

The value of 1/root 10 when root 10 = 3.162 ?

OpenStudy (anonymous):

Please help ! Exam on Monday !

OpenStudy (anonymous):

1/root 10 =( root 10 )/10 and so on..

OpenStudy (anonymous):

What is root 10 / 10 ?

OpenStudy (anonymous):

( root 10 )/10 = 3.162 /10 = 0.3162

OpenStudy (anonymous):

Tnx a loot !